OpenStudy (anonymous):

HOW DO YOU MULTIPLY 1/2 BY 116?????

OpenStudy (anonymous):

You can turn 116 into 116/1, then multiply 1/2 X 116/1

OpenStudy (anonymous):

you change 1/2 into a decimal which is 0.5 and multiply that by 116 which gives you 58

OpenStudy (anonymous):

That works too :) @caligirl09

OpenStudy (anonymous):

Thanks 4 the medal @bibby

OpenStudy (anonymous):

how do you change it into a decinal?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

you divide 1 by 2

OpenStudy (studygurl14):

\(\large\frac{1}{2}\times116\rightarrow\frac{1}{2}\times\frac{116}{1}\rightarrow\frac{116}{2}\) What is 116/2? @KierseyClemons

OpenStudy (anonymous):

@KierseyClemons

OpenStudy (anonymous):

2 divded by 166?

OpenStudy (studygurl14):

No...166 divided by 2

OpenStudy (anonymous):

I mean 16

OpenStudy (studygurl14):

Another way to do it is to change 1/2 to 0.5 and multiply it by 116. Just so you know.

OpenStudy (studygurl14):

and no. not 16

OpenStudy (anonymous):

no you change 1/2 into into 0.5 by dividing 1 by 2 and then you multiply 166 by 0.5

OpenStudy (anonymous):

and get 58 which is your answer

OpenStudy (studygurl14):

@KierseyClemons What is 116 divided by 2?

OpenStudy (studygurl14):

Yea @caligirl09 it is 58. @KierseyClemons

OpenStudy (anonymous):

did youmidactually workingitout soI cancopy itfornotes?i getconfused @StudyGurl14

OpenStudy (anonymous):

116*1/2 1/2=0.5 (1 divided by 2) 0.5*116=58

OpenStudy (anonymous):

like the setup and everything @StudyGurl14

OpenStudy (anonymous):

Is what I wrote what you needed @KierseyClemons

OpenStudy (anonymous):

like I mean work itout..like carrying numbers ect

OpenStudy (anonymous):

ialways get the wrong answer idk why

OpenStudy (studygurl14):

I don't get what you're looking for...look up there ^ for my steps.